# High Dimensional Distributed Gradient Descent with Arbitrary Number of Byzantine Attackers

### v2

@@ Abstract
Robust distributed learning with Byzantine failures has attracted extensive research interests in recent years. However, most of existing methods suffer from curse of dimensionality, which is increasingly serious with the growing complexity of modern machine learning models. In this paper, we design a new method that is suitable for high dimensional problems, under arbitrary number of Byzantine attackers. The core of our design is a direct high dimensional semi-verified mean estimation method. Our idea is to identify a subspace first. The components of mean value perpendicular to this subspace can be estimated via gradient vectors uploaded from worker machines, while the components within this subspace are estimated using auxiliary dataset. We then use our new method as the aggregator of distributed learning problems. Our theoretical analysis shows that the new method has minimax optimal statistical rates. In particular, the dependence on dimensionality is significantly improved compared with previous works.
1 Introduction
@@
Many modern machine learning tasks require distributed computing and storage. In such systems, there are many worker machines, which are coordinated by a centralized server, called the master machine. Each worker machine has a local training dataset which is never uploaded to the master. Instead, worker machines only need to compute an update to the current model, and send it to the master. Such setting is called Federated Learning (FL) [1], which has drawn considerable attention in recent years [2, 3]. FL has advantages in both computational efficiency and privacy preserving. With the rapid growth of the amount of data, storing and computing data in a distributed manner becomes necessary. Moreover, in many sensitive applications, such as healthcare, recommendation systems and self-driving cars, FL can protect users’ privacy, since the model is trained without direct access to local datasets stored in clients.
Despite significant advantages and widespread applications, FL is still facing many challenges [3]. One critical issue is that not all machines are trustable [4]. In large scale FL problems, some of worker machines may send wrong information to the master due to various reasons, including system crashes, faulty hardware, communication errors, and even malicious attacks [5, 6, 7, 8, 9, 10]. Unfortunately, the vanilla FL [1] is highly susceptible to these corrupted worker machines. Without proper handling, the trained model weights may be far from optimal. Consequently, evaluating the impacts of these corrupted worker machines, and designing corresponding defense strategies, have become critical tasks in the research community.
Abnormal behaviors of worker machines are usually unpredictable. To build a robust system that is resilient to all corruptions, it is necessary to consider the worst case. With such intuition, researchers typically model these faults as Byzantine failure [11]. Under this model, for some worker machines, the messages sent to the master are manipulated carefully by a malicious adversary. It is assumed that the adversary knows the underlying distribution of data, as well as the learning algorithm. With these information, the adversary designs an optimal strategy to deteriorate the model performance as much as possible. Therefore, Byzantine failure is the most harmful among all corruptions. If our learning algorithm can defend against Byzantine attack, it should also be robust under other possible failures.
There have been extensive research on Byzantine robust distributed learning. Suppose that there are $m$ worker machines, with each machine storing $n$ samples. The attacker can manipulate up to $q$ worker machines. Most of previous works focus on the case with $q<m/2$, i.e. less than half of worker machines are attacked, such as Krum [12], geometric median of mean [13], coordinate-wise median and trimmed mean [14]. There are also some previous results on the case with arbitrary number of attackers, which allows $q>m/2$. Actually, without additional information, it is impossible to learn the model precisely. Therefore, these works typically rely on the help of an additional clean dataset, so that the real gradient can be estimated by combining the gradient vectors uploaded by these worker machines and the information from the auxiliary clean samples. This line of works include [15, 16, 17, 18], see Section 2 for further discussion.
Nevertheless, a common problem for all these methods is the curse of dimensionality. The final statistical errors grow with the dimensionality $d$, which is the number of model parameters. As a result, these methods are not suitable in modern large scale machine learning models. A progress towards high dimensional robust federated problem with $q<m/2$ was made in [19], which is based on some high dimensional robust mean estimators proposed in recent years [20, 21, 22, 23, 24, 25, 26, 27]. These new estimators are statistical and computational efficient. According to [19], with these robust mean estimators as gradient aggregators in Byzantine robust federated learning tasks, if $q<m/2$, then the convergence rate of learning error is nearly minimax optimal. However, to the best of knowledge, high dimensional cases with $q\geq m/2$ has not been solved before.
In this paper, we propose a new approach to distributed gradient descent in high dimensions, which allows arbitrary number of Byzantine attackers. While the focus is primarily on the case with $q\geq m/2$, our method also works with $q<m/2$. Similar to previous works [15, 16, 17, 18], we assume that an auxiliary dataset is available. The main improvements compared with [15, 16, 17, 18] is that our method can overcome the curse of dimensionality. The detailed steps are shown as follows.
Firstly, we propose a direct semi-verified mean estimation method. Semi verified mean estimation is the problem of estimating the mean of a distribution, by combining the information of a small clean dataset and a large untrusted dataset, in which a large fraction of samples may be corrupted [28]. Semi verified mean estimation algorithm is important since it can serve as the aggregator function in distributed learning. [28] shows that semi-verified mean estimation can be converted to list-decodable mean estimation problem, which aims at generating a list of $K$ hypotheses $\hat{\mu}_{1},\ldots,\hat{\mu}_{K}$ such that one of them is close to the ground truth $\mu^{*}$. With a hypotheses list of length $K$, we can then use $O(\ln K)$ clean samples to conduct a hypothesis testing to pick one from the hypotheses list. However, we argue that the indirect semi-verified mean estimation method by conversion to list-decodable mean estimation is a bit complex and sample inefficient. Therefore, we propose a direct method to semi-verified mean estimation. The intuitive idea of our approach can be explained as follows. The first step is to filter out some samples and identify a subspace in which the components of $\mu^{*}$ is hard to estimate using the corrupted data. These components are then estimated using auxiliary clean dataset. In directions that are perpendicular to the subspace, we just use the projected sample mean of the corrupted dataset. Our design is partially motivated by a list-decodable mean estimation method in [29].
Secondly, we conduct a theoretical analysis of our new proposed semi-verified mean estimation algorithm. Under the condition that the covariance matrix is bounded by $\sigma^{2}\mathbf{I}$, and the fraction of good samples among the untrusted dataset is at least $\alpha$, we obtain an upper bound of the estimation error with respect to $\sigma$ and $\alpha$. Our analysis considers two models separately. The first one is additive contamination model. Under this model, attacked samples are injected into a set of good samples. As a result, the distribution of good samples is not influenced by the adversary. Another way to understand this model is that the adversary randomly pick $1-\alpha$ fraction of samples from the corrupted dataset and change their values arbitrarily, The second one is strong contamination model, under which the adversary can pick $1-\alpha$ fraction of samples arbitrarily instead of randomly, and then modify their values to maximize their negative impacts on the estimation performance. As a result, the distribution of good samples is also altered after contamination. Compared with additive contamination model, the adversary is more powerful under strong contamination model. Consequently, the theoretical bound under strong contamination model is also larger. Apart from the upper bound of our new method, we also obtain the information theoretic minimax lower bound under both additive and strong contamination model. The result shows that our upper and lower bounds match, indicating that our method is optimal, in the sense that the sample efficiency can not be further improved in general.
Finally, we apply the semi-verified mean estimation algorithm into Byzantine robust distributed learning problem, with the gradient vector returned by each worker machine being regarded as a sample. We conduct theoretical analysis of the performance of distributed learning under Byzantine attack.
The remainder of this paper is organized as follows. In section 2, we provide a review of previous literatures. Section 3 shows our semi-verified mean estimation method. Our theoretical analysis is shown in Section 4. After that, Section 5 discusses the distributed learning algorithm. Finally, we offer numerical results and concluding remarks in Section 6 and Section 7, respectively. Proofs are shown in the appendix.
2 Related Work
2.1 Robust mean estimation
Distributed learning relies on accurate estimation of the statistical mean of gradients gathered from worker machines. Therefore, we provide a brief review of robust mean estimation first. Estimation with corrupted samples has been an important task in statistics since 1960s [30, 31]. Traditional robust estimation methods such as trimmed mean behaves badly as dimensionality increases. An exception is Tukey median [32], which is sample efficient in high dimensional spaces. However, Tukey median is not practical, since the computation can not finish in polynomial time. [33, 20, 21] proposed convex programming and filtering methods for high dimensional robust mean estimation. Since then, a flurry of works has emerged, focusing on improving the time complexity [34], extending to sparse problems [35, 22], and improving the performance under stronger tail assumption with sum of squares method [36]. These methods works if the proportion of contaminated samples is less than $1/2$.
If majority of samples are corrupted, then it is impossible to estimate the mean with a single solution. Suppose that the fraction of clean samples is ensured to be at least $\alpha$, i.e. $q/m<1-\alpha$, then the adversary can produce $\Omega(1/\alpha)$ clusters of points, with one of them composed of clean samples, while others are corrupted. In this case, there is no way to determine which cluster is correct. Therefore, in recent works, the goal of mean estimation is relaxed to generating a list of hypotheses, with the guarantee that at least one of the hypotheses is close to the ground truth $\mu^{*}$. This paradigm is called list-decodable learning. Typically, the length of generated hypotheses list need to be controlled within $O(1/\alpha)$. We hope to make the minimum estimation error among the hypotheses list to be as low as possible. [28] proposed an SDP-based method, achieving error of $O(\sigma\sqrt{\ln(1/\alpha)/\alpha})$. For spherical Gaussians, a multi filter method can achieve a better rate [37]. Several recent works improves the time complexity of list-decodable mean estimation. [29] proposed SIFT (Subspace Isotropic Filtering), which achieves nearly PCA time complexity. [38] introduced an optimization based approach with linear time complexity.
Apart from the untrusted data, if we also have some auxiliary clean samples, then it is possible to determine one estimated value from the hypotheses list, which is close to the ground truth with high probability. [28] named it semi-verified learning, and claims that with $K$ hypotheses, one needs $O(\ln K)$ clean samples to determine the correct one.
2.2 Federated Learning with less than half Byzantine machines
Suppose that there are $m$ worker machines, among which $q$ of them being attacked. Each working machine has $n$ samples. The total size of dataset is $N=mn$. Here we briefly review the previous works with $q<m/2$.
[12] proposed Krum, which collects gradient vectors from these $m$ worker machines, and then select a vector that is closest to its $m-q$ neighbors. It is expected that the selected one is close to the ground truth. However, the error bound is suboptimal. [13] uses a geometric median of mean method, with a suboptimal error rate of $\sqrt{qd/N}$. [14] made the first step towards optimal statistical rates of Byzantine robust federated learning. Two methods are proposed: coordinate-wise median and coordinate-wise trimmed mean. Under the assumption that the gradient has bounded skewness, the coordinate-wise median method achieves $\tilde{O}(q\sqrt{d}/(m\sqrt{n})+d/\sqrt{nm}+\sqrt{d}/n)$. If $n\gtrsim m$, then this method is optimal in $q$, $n$ and $m$. Under the subexponential tail assumption, the coordinate-wise trimmed mean achieves $\tilde{O}(d(q/(m\sqrt{n})+1/\sqrt{nm}))$, which is also optimal. A new approach based on Huber loss minimization was proposed in [39], which has advantages of not requiring precise knowledge of $q$ and suitability to worker machines with unequal number of samples. However, these methods are not optimal in $d$. As a result, in large scale models, the performance of these models are not desirable. There are also several works on non-i.i.d samples, such as RSA [40] and RAGA [41].
Recently, [19] shows that by using several recent high dimensional robust mean estimation methods [20, 21, 22, 23, 24, 25, 26, 27] as the gradient aggregator functions, the error rate of federated learning becomes nearly optimal not only in $m$, $n$, $q$, but also in dimensionality $d$, up to a logarithm factor.
2.3 Federated learning with arbitrary number of Byzantine machines
The condition that less than half worker machines are attacked may not hold in practice. In this case, the methods mentioned above no longer works. If $q>m/2$, without additional information, it is impossible to provide a precise solution. [15] provides a simple method with the help of a clean auxiliary dataset. The basic idea is to calculate the gradient vector only using the auxiliary dataset as a baseline, and then filter out gradient vectors uploaded from worker machines that are too far away from such baseline. A slightly different method was proposed in [16], which introduces reputation scores as the weights for gradient. However, the error rate of these methods is $\|\hat{\mathbf{w}}-\mathbf{w}^{*}\|=O(\sqrt{d/N_{A}})$, with $N_{A}$ being the size of auxiliary dataset. This rate is not significantly improved compared with learning with auxiliary data only, and is undesirable in high dimension problems. A significantly different method called Zeno was proposed in [17], which formulates a stochastic descent score to filter out adversarial workers. [18] proposed Zeno++, an improvement to Zeno. The main advantage of Zeno++ is the suitability under asynchronous setting. Moreover, by computing a first order approximation of the gradient descent score, this method significantly improves the computational efficiency. However, the error bound is also significantly suboptimal. In particular, the error rate of Zeno and Zeno++ is still $O(\sqrt{d/N_{A}})$ (this comes from calculating the term $V_{3}$ in Theorem 1 in [18]), which also scales badly with the increase of dimensionality. To the best of our knowledge, our work is the first attempt to solve Byzantine tolerant distributed learning problem in high dimensions.
3 Semi Verified Mean Estimation
In this section, we analyze the semi-verified mean estimation problem, in which we hope to estimate the mean of distribution by combining a small set of clean samples and a large set of corrupted samples.
To begin with, denote $S_{0}$ as the set of all untrusted samples, $\mathcal{B}$ as the set of attacked samples. Suppose that the ratio of clean samples is at least $\alpha$, then $|\mathcal{B}|\leq(1-\alpha)N$. $S_{0}^{*}:=S_{0}\setminus\mathcal{B}$ is the set of clean samples. Denote $\mu^{*}=\mathbb{E}[\mathbf{X}]$ and $V^{*}=\mathbb{E}[(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}]$ as the statistical mean and covariance matrix, respectively, in which $\mathbf{X}\sim D$. The samples before contamination are denoted as $\mathbf{X}_{1},\ldots,\mathbf{X}_{N}$, which are not observed. Instead, we can observe a corrupted one $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$, in which
@@ 3 Semi Verified Mean Estimation
$\mathbf{Y}_{i}=\left\{\begin{array}[]{ccc}X_{i}&\text{if}&i\notin\mathcal{B}\\ \star&\text{if}&i\in\mathcal{B},\end{array}\right.$ (3)
in which $\star$ means arbitrary value determined by the attacker. Apart from the corrupted dataset, an auxiliary clean set $A$ of clean samples is available, with size $N_{A}=|A|$. We mainly consider the case such that $N_{A}$ is small.
Here show the definition of two models, i.e. additive contamination model and the strong version. The definitions are from [20, 22]. We would like to remark that if less than half of samples are attacked, then the difference between these two models is not significant. However, our result indicate that if majority of samples are attacked, the parameter selection and error bounds are significantly different between these two models, especially if $\alpha$ is small.
@@
(Additive Contamination Model) Given a parameter $\alpha$ and a distribution family $\mathcal{F}$ on $\mathbb{R}^{d}$, suppose that there are $N$ samples, in which at least $\alpha N$ of them are drawn from an unknown distribution $D\in\mathcal{F}$. The values of other samples are determined arbitrarily by the attacker.
@@
(Strong Contamination Model) Given a parameter $\alpha$ and a distribution family $\mathcal{F}$ on $\mathbb{R}^{d}$, suppose that there are $N$ samples drawn from an unknown distribution $D\in\mathcal{F}$. The adversary is allowed to inspect the samples, and change the values of at most $(1-\alpha)N$ samples arbitrarily.
@@
Under additive contamination model, samples in $S_{0}^{*}$ follow distribution $D$. Therefore, an equivalent formulation of additive contamination model is that suppose $N$ samples are all generated from $D$, and then the attacker selects at most $(1-\alpha)N$ samples randomly, and alter their values arbitrarily. Strong contamination model is crucially different, since the adversary can pick samples carefully instead of randomly, before altering their values. As a result, the distribution of remaining clean samples can be far away from $D$. Therefore, strong contamination model is more challenging. Even if we can identify good samples from the corrupted dataset, estimation of $\mu^{*}$ is still not guaranteed. Practical scenario usually lie in between these two models. In distributed learning problems, the attacked working machines are usually not randomly selected, thus additive contamination model underestimates the impact of Byzantine attacks. On the contrary, strong contamination model tend to overestimate such impact, since it is unlikely for the adversary to gather full information and design optimal attack strategy. In this work, we analyze both two models.
Now we clarify the knowledge of the adversary.
Assumption 1.
The adversary knows the underlying distribution $D$. Moreover, under strong contamination model, the adversary also knows original sample values $\mathbf{X}_{1},\ldots,\mathbf{X}_{N}$. However, the adversary has no knowledge of the clean auxiliary data $A$.
Assumption 1 requires that the adversary has no knowledge of $A$. This is both practical and necessary. In real scenarios, it is usually not hard to collect only a small set of auxiliary clean data from a reliable source that is not inspected by the attacker, thus it is practical to assume that adversary does not know sample values in $A$. To see why Assumption 1 is necessary, consider a simple attack strategy: if the adversary knows $A$, just flip half of clean samples with respect to the sample mean of $A$. Recall that the unobserved clean dataset is denoted as $\{\mathbf{X}_{1},\ldots,\mathbf{X}_{N}\}$. The attacker can just let $\mathbf{Y}_{i}=2\mu(A)-\mathbf{X}_{i}$ for half of samples, in which $\mu(A)$ is the sample mean of auxiliary dataset $A$, and $\mathbf{Y}_{i}=\mathbf{X}_{i}$ for others. As a result, the corrupted set will have two clusters that are placed symmetrically around $\mu(A)$, and thus it is impossible to distinguish which one is closer to the ground truth $\mu^{*}$. Therefore, it is necessary to assume that the adversary can not observe auxiliary dataset $A$, otherwise the overall estimation accuracy can not be better than using $A$ only.
Here we describe the algorithm as following, and the formal statement is shown in Algorithm 1. This algorithm is partially motivated by [29]. Initialize $S=S_{0}$. Firstly, we use a prefilter (step 1) to remove the samples whose norms are too large. The prefilter is practically not necessary, and the threshold is also not important. This step is mainly used to simplify theoretical analysis. Then we remove points in $S$ iteratively. We manage to ensure that most of the removed points are corrupted, so that $S$ gradually becomes cleaner.
In each step, we first calculate the sample mean and covariance:

### v3

@@ Abstract
Adversarial attacks pose a major challenge to distributed learning systems, prompting the development of numerous robust learning methods. However, most existing approaches suffer from the curse of dimensionality, i.e. the error increases with the number of model parameters. In this paper, we make a progress towards high dimensional problems, under arbitrary number of Byzantine attackers. The cornerstone of our design is a direct high dimensional semi-verified mean estimation method. The idea is to identify a subspace with large variance. The components of the mean value perpendicular to this subspace are estimated using corrupted gradient vectors uploaded from worker machines, while the components within this subspace are estimated using auxiliary dataset. As a result, a combination of large corrupted dataset and small clean dataset yields significantly better performance than using them separately. We then apply this method as the aggregator for distributed learning problems. The theoretical analysis shows that compared with existing solutions, our method gets rid of $\sqrt{d}$ dependence on the dimensionality, and achieves minimax optimal statistical rates. Numerical results validate our theory as well as the effectiveness of the proposed method.
Introduction
@@
Many modern machine learning tasks require distributed computing and storage. In such systems, there are many worker machines, which operate under the coordination of a centralized server, known as the master machine. During the whole training process, worker machines only need to compute an update to the current model, and send it to the master, while the local dataset remains private. This framework is called Federated Learning (FL) (mcmahan2017communication), which has drawn considerable attention in recent years (zhang2021survey; kairouz2021advances). With advantages in both computational efficiency and privacy protection, FL has been widely applied in various areas, including smart devices, industrial engineering and healthcare (li2020review; li2020federated).
Despite significant advantages and widespread applications, FL is still facing many challenges (kairouz2021advances). One critical issue is that not all machines are trustable (lyu2020threats). In large scale FL systems, some worker machines may send wrong gradients to the master due to various reasons, including system crashes, faulty hardware, communication errors, and even malicious attacks (bagdasaryan2020backdoor; bhagoji2019analyzing; fang2020local; sun2021data; luo2021feature). These faults are typically modeled as Byzantine failure (lamport1982byzantine), such that some workers are manipulated by an adversary, and send wrong gradient vectors to the master.
There have been extensive research on Byzantine robust distributed learning. Many popular methods (blanchard2017machine; guerraoui2018hidden; chen2017distributed; yin2018byzantine) have two issues that make them not perfect. Firstly, these methods only allow less than half worker machines to be Byzantine. Secondly, the statistical risk grows with dimension $d$, which is serious in modern large scale learning models. For the former issue, several new methods suitable for arbitrary number of Byzantine attackers have been proposed recently (cao2019distributed; regatti2020bygars; xie2019zeno; xie2020zeno++; cao2020fltrust). These methods rely on the help of a small auxiliary clean dataset, which is necessary since the breakdown point of robust statistics can not exceed $1/2$ (hampel1971general). The latter problem, i.e. curse of dimensionality, has been addressed by some recent works based on high dimensional robust statistics (diakonikolas2023algorithmic), such as (su2018securing; shejwalkar2021manipulating; zhu2022c), which have constant dependence on $d$. However, to the best of our knowledge, there are no previous methods that can solve both two issues mentioned above. Unfortunately, in modern federated learning problems, it may happen that models have large scale and most workers are not trusted.
In this paper, we tackle both problems mentioned above simultaneously. In particular, we propose a novel approach to high dimensional distributed learning under arbitrary number of Byzantine attackers, with the help of a small auxiliary dataset. The core of our new approach is a new algorithm for semi-verified mean estimation, whose goal is to estimate the statistical mean by combining a small clean dataset and a large untrusted dataset (charikar2017learning). The basic idea is to filter out some gradient vectors that are very likely to come from Byzantine workers, and then identify a linear subspace, in which the components in this subspace are hard to estimate using these corrupted gradient vectors. These components are then estimated using auxiliary clean dataset. For other components that are perpendicular to this subspace, we just use the projected sample average. Our design is partially motivated by (diakonikolas2021list). A theoretical analysis is provided, including the upper bound of our algorithm and the information-theoretic minimax lower bound. The new semi-verified mean estimation algorithm is then used as the gradient aggregator for robust distributed learning problem.
Preliminaries
Problem Statement of Byzantine Robust Distributed Learning
To begin with, we formalize the problem of distributed learning under Byzantine failures. Our formalization follows (chen2017distributed; yin2018byzantine; zhu2023byzantine).
Suppose there is a master machine $W_{0}$ and $m$ worker machines, $W_{1},\ldots,W_{m}$. A clean dataset $A=\{\mathbf{Z}_{01},\ldots,\mathbf{Z}_{0N_{A}}\}$ is stored in the master, with size $|A|=N_{A}$. Each worker machine $W_{i}$ has $n$ samples $\mathbf{Z}_{i1},\ldots,\mathbf{Z}_{in}$. All samples stored in both master and worker machines are identically and independently distributed (i.i.d), following a common distribution $Q$.
Let $f(\mathbf{w},\mathbf{z})$ be a loss function of a parameter vector $\mathbf{w}\in\Omega\subseteq\mathbb{R}^{d}$, in which $\Omega$ is the parameter space. The population risk function is defined as
$F(\mathbf{w})=\mathbb{E}[f(\mathbf{w},\mathbf{Z})],$ (1)
in which $\mathbf{Z}\sim Q$. Our goal is to learn the minimizer of $F$, i.e. $\mathbf{w}^{*}=\underset{\mathbf{w}\in\Omega}{\arg\min}F(\mathbf{w})$.
The FL framework runs as following. In each iteration, the master machine broadcasts current parameter $\mathbf{w}_{t}$ to all worker machines. Then these worker machines calculate the gradient using the local dataset:
$\mathbf{X}_{i}(t)=\frac{1}{n}\sum_{j=1}^{n}\nabla f(\mathbf{w}_{t},\mathbf{Z}_{ij}).$ (2)
For any benign workers, the master can receive $\mathbf{X}_{i}(t)$ exactly. On the contrary, Byzantine machines can send arbitrary messages determined by the attacker. Denote $\mathbf{Y}_{i}(t)\in\mathbb{R}^{d}$ as the vector received by the master at $t$-th iteration, then
@@
$\mathbf{Y}_{i}(t)=\left\{\begin{array}[]{ccc}\mathbf{X}_{i}(t)&\text{if}&i\notin\mathcal{B}\\ \star&\text{if}&i\in\mathcal{B}.\end{array}\right$. (5)
Furthermore, we can calculate the gradient using the clean dataset stored in the master that is never corrupted, i.e.
$\mathbf{X}_{0}(t)=\frac{1}{N_{A}}\sum_{j=1}^{N_{A}}\nabla f(\mathbf{w}_{t},\mathbf{Z}_{0j}).$ (6)
Algorithm 1 Robust Distributed Gradient Descent
Input: Master machine $W_{0}$, worker machines $W_{1},\ldots,W_{m}$
Parameter:Initial weight parameter $\mathbf{w}_{0}\in\Omega$, step length $\eta$, total number of iterations $T$
Output:Estimated weight $\hat{\mathbf{w}}$
1:for $t=0,1,\ldots,T-1$ do
2:  Master machine: broadcast current parameter $\mathbf{w}_{t}$ to all worker machines;
3:  for $i\in[m]$ in parallel do
4:   Worker machine $i$: compute local gradient $\mathbf{X}_{i}(t)$ using (2);
5:   if $i$ is normal machine then
6:    send $\mathbf{X}_{i}(t)$ to master;
7:   else
8:    send arbitrary $d$ dimensional vector to master;
9:   end if
10:  end for
11:  Master machine: Receive $\mathbf{Y}_{i}(t)$, $i=1,\ldots,m$ from each worker machine;
12:  Calculate aggregated gradient $g(\mathbf{w}_{t})$ using $\mathbf{Y}_{i}(t)$, $i=1,\ldots,m$ and clean dataset in $W_{0}$;
13:  Update parameter $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta g(\mathbf{w}_{t})$ ;
14:end for
We would like to clarify that the adversary knows our algorithm, as well as $m$ gradient vectors $\mathbf{X}_{i}(t)$, $i=1,\ldots,m$. However, the adversary has no knowledge of the clean dataset $A$. This is practical since it is usually not hard to collect only a small set of auxiliary clean data from a reliable source that is not inspected by the attacker.
It remains to discuss how to calculate the aggregated gradient $g(\mathbf{w}_{t})$. This problem can be formulated as Semi-verified mean estimation (see Definition 1), which was first defined in (charikar2017learning).
Problem Statement of Semi-verified Mean Estimation
The aggregator $g(\mathbf{w}_{t})$ estimates $\nabla F(\mathbf{w}_{t})$ using $\mathbf{Y}_{i}(t)$, $i=1,\ldots,m$ as well as $\mathbf{X}_{0}(t)$, which is calculated using (6). Among $m$ worker machines, $\alpha m$ of them are ensured to be benign, while others are probably Byzantine. Therefore, we state semi-verified mean estimation problem as follows.
@@
(Semi-verified mean estimation problem) Suppose $\mathbf{X}_{1},\ldots,\mathbf{X}_{m}$ are i.i.d with mean $\mu^{*}$ and covariance matrix $V^{*}$. $\mathcal{B}\subset S_{0}=[m]$ is the set of attacked samples. If $i\notin\mathcal{B}$, then $\mathbf{Y}_{i}=\mathbf{X}_{i}$, otherwise $\mathbf{Y}_{i}$ is an arbitrary vector determined by the attacker. In addition, we have a sample $\mathbf{X}_{0}$ with mean $\mu^{*}$ and covariance matrix $V_{A}^{*}$, which is guaranteed to be unattacked. The task is to estimate $\mu^{*}$ using untrusted samples $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m}$ and the clean sample $\mathbf{X}_{0}$.
Consider that the semi-verified mean estimation need to be repeated for all iterations, we omit timestep $t$ in Definition 1. Each sample $\mathbf{Y}_{i}$ corresponds to the gradient vector from a worker machine $W_{i}$. At iteration $t$, $\mu^{*}=\nabla F(\mathbf{w}_{t})$. Denote $V_{0}$ as the covariance matrix of $\nabla f(\mathbf{w}_{t},\mathbf{Z}_{ij})$, then according to (2) and (6), $V^{*}$ and $V_{A}^{*}$ in Definition 1 become
$V^{*}=\frac{V_{0}}{n},V_{A}^{*}=\frac{V_{0}}{N_{A}}.$ (7)
Finally, we discuss two attack models that are different on whether the attacked samples are randomly (Definition 2) or carefully selected (Definition 3).
@@
(Additive Contamination Model) The adversary randomly pick at most $(1-\alpha)m$ samples from $\mathbf{X}_{1},\ldots,\mathbf{X}_{m}$, and alter their values arbitrarily.
Definition 3.
(Strong Contamination Model) The adversary picks $(1-\alpha)m$ samples from $\mathbf{X}_{1},\ldots,\mathbf{X}_{m}$ according to their values, and alter their values arbitrarily.
These definitions follow (diakonikolas2016robust; diakonikolas2020outlier). Under additive contamination model, adversarial samples are randomly selected, so the distribution of remaining clean samples remains unchanged; equivalently, there are already $\alpha m$ samples, and the adversary injects the remaining $(1-\alpha)m$ corrupted samples with arbitrary values. Under strong contamination model, since adversarial samples are carefully selected by the attacker, and thus those remaining samples follow a different distribution. Even if good samples can be identified, estimation of $\mu^{*}$ is still not guaranteed.
@@
Practical scenarios usually lie in between these two models. In distributed learning problems, the attacked worker machines are usually not randomly selected, thus additive contamination model underestimates the impact of Byzantine attacks. On the contrary, strong contamination model tends to overestimate such impact, since it is unlikely for the adversary to gather full information and design optimal attack strategy. In this work, we analyze both two models.
Semi-verified Mean Estimation
Our algorithm for semi-verified mean estimation is shown in Algorithm 2, which is partially motivated by the Subspace Isotropic Filtering (SIFT) algorithm proposed in (diakonikolas2021list). The idea is illustrated in Figure 1. The procedures are explained as follows.
Algorithm 2 Semi verified mean estimation
Input: Untrusted samples $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m}$ (from (5)), and clean sample $\mathbf{X}_{0}$ (from (6))
Parameter: $p$, $\lambda_{c}$
Output: Estimated mean $\hat{\mu}$
1: Initialize $S=S_{0}$;
[…]
3:while True do
4:  Calculate sample mean $\mu(S)$ and sample covariance $V(S)$ using (8) and (9);
5:  Conduct spectral decomposition of $V(S)$, such that $V(S)=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$, in which $\mathbf{U}$ is orthogonal, and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})$, $\lambda_{1}\geq\ldots\geq\lambda_{d}$;
6:  if $\lambda_{p}\geq\lambda_{c}$ then
7:   Let $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}$, in which $\mathbf{U}_{p}\in\mathbb{R}^{d\times p}$ is the first $p$ columns of $\mathbf{U}$;
8:   For each $i\in S$, calculate $\tau_{i}$ using (10);
9:   $\tau_{\max}=\max_{i\in S}\tau_{i}$;
10:   For each $i\in S$, remove $i$ from $S$ with probability $\tau_{i}/\tau_{\max}$;
11:  else
12:   break;
13:  end if
14:end while
15: Calculate $V(S)$, $\mathbf{P}$;
[…]
17:return $\hat{\mu}$;
(a) Initialization and prefiltering (Step 1-2).
(b) Iterative filtering (Step 3-14).
(c) Calculate sample mean and projection matrix (Step 15).
(d) Calculate final estimate $\hat{\mu}$ (Step 17).
Figure 1: A two-dimensional illustration of the semi-verified mean estimation method shown in Algorithm 2. $\mu^{*}$ is represented by the red triangle. Benign and attacked samples correspond to green and black dots, respectively. The orange plus sign denotes $\mathbf{X}_{0}$.
(1) Initialization and prefiltering. Let $S_{0}=\{1,\ldots,m\}$ be the indices of all untrusted samples and initialize $S=S_{0}$. We use a prefilter (step 2) to remove the samples whose norms are too large. This step is mainly used to simplify the theoretical analysis, and may not be necessary in practice.
(2) Iterative filtering. The algorithm then sanitizes the dataset by removing some samples that are highly likely to be attacked. In each iteration, sample mean and covariance matrix are calculated first: